OpenStudy (sloppycanada):

Check my work? Central Angles Find the measure of the central angle with a radius of 13 inches and the area of a sector measuring 42.25pi square inches.

OpenStudy (sloppycanada):

\[\frac{ 7605 }{ \frac{ 1 }{ 2 } (13^2)}\] I got this because I translated 42.25pi to degrees (7605)

OpenStudy (sloppycanada):

From there, it should be 90 degrees for the central angle.

OpenStudy (sloppycanada):

@TheSmartOne If you have time?

TheSmartOne (thesmartone):

\(\sf\Large Sector~Area = \frac{\theta}{2}\cdotr^2\) Plugging in the info given: \(\sf\Large 42.25\pi = \frac{\theta}{2}\cdot(13)^2\) Solve for \(\theta\)

OpenStudy (sloppycanada):

I thought sector area was... \[A = \frac{ 1 }{ 2 }r^2\theta\]

OpenStudy (sloppycanada):

If I follow your equation, I'm looking at... 1.57

OpenStudy (sloppycanada):

Which doesn't seem right.

TheSmartOne (thesmartone):

\(\sf\Large Sector~Area = \frac{\theta}{2}\cdot r^2\) Plugging in the info given: \(\sf\Large 42.25\pi = \frac{\theta}{2}\cdot(13)^2\) Solve for \(\theta\)

TheSmartOne (thesmartone):

The formula you stated is the same.

TheSmartOne (thesmartone):

theta/2 * r^2 = 1/2 * theta * r^2 same thing :P

OpenStudy (sloppycanada):

\[84.5\pi=\theta*169\]

OpenStudy (sloppycanada):

\[84.5\pi/169 = \theta\]

OpenStudy (sloppycanada):

I'm still getting 1.57

TheSmartOne (thesmartone):

why are you making it a decimal?

TheSmartOne (thesmartone):

84.5/169 = 845/1690 = 1/2

OpenStudy (sloppycanada):

It's what my calculator says.

OpenStudy (sloppycanada):

Alright. But 1/2 isn't in degrees. and 50 degrees isn't one of my choices.

OpenStudy (sloppycanada):

@Astrophysics if you get the chance.

TheSmartOne (thesmartone):

It's 1/2 pi 0.5* pi

TheSmartOne (thesmartone):

and that's in radians ^